I found a strange jump of value in my data taken with tdsdata.
I couldn't find same jump in a playback of DataViewer, so I think this is a problem of tdsdata.
Be careful when you use tdsdata!
The attached file is an example of jumped data.
I try to get data with allegra and op440m, and both has same kind of jump.
(A downsampling or interpolation may be wrong.)
Rana said there is a fixed version of tdsdata in some PC, but 64bit linux may not have.
I try it tomorrow.
While continuing our efforts to lock, we noticed the procedure failed at a point it had gotten past last night: turning on the bounce/roll filters in MICH, PRC, and SRC. We checked the MICH transfer function and noticed that the unity gain point was ~10 Hz, well below the bounce modes. We tried increasing the gain but found saturation, and Rob suggested that there could be misalignment on the AP table, which Steve worked on today. We went out and found two of the PDs (ASDD133 and AS166) to be badly misaligned probably due to a bumped optic upstream. We re-aligned.
I checked a broken QPD, which was placed for PSL angle monitor, and finally I cocluded one segment of the quadrant diode was broken.
The broken segment has a offset voltage of -0.7V after 1st I-V amplifier. It means the diode segment has a current offset without any injection of light.
Tomorrow I will check a new QPD for replacement.
The control room video is showing us a false ETMY image. Who worked on the ETMY camera or video today??!!
The attached plot shows that someone broke the MC_SUM_MON channel around 10:30 AM this past Wednesday the 11th. This is the EPICS monitor of the MC error point.
Come forward now with your confession and I promise that I won't let Steve hurt you.
We centerd the input of WFS QPD.
c1susvme1 is behaving weirdly. I've restarted it several times but its computation time is hanging out around 260 usec, making it useless for suspension control and locking. I also found a PS/2 keyboard plugged in, which doesn't work, so I unplugged it. It needs to be plugged into a PS/2 keyboard/mouse Y-splitter cable.
I have been getting ready for the annual safety inspection in the past 2-3 days.
I added a PA current limiter.
It is only a voltage devider (composed with 3.09k and 1.02k resiste) between DAC and PA current adjustment input.
The output range of DAC is +/- 10[V] and the conversion factor of PA current adjustment is 0.84[A/V] (measured value), so the PA current adjustment is limited +/- 2.1[A] ( 10[V]*1.02k/(1.02k+3.09k)*0.84[A/V] ).
Actually, the manual of the PA tells that the conversion factor is 0.25[A/V].
There is 3 possibility.
1) There are some mistakes in channels of digital system.
2) The PA manual is wrong.
2-1) The conversion factor of current adjustment is wrong.
2-2) The conversion factor of current monitor is wrong.
I measured the signal of current adjustment and current monitor directly, and confirm that they are consistent to the value monitord from MEDM.
Hence the PA manual must be wrong, but I don't know which factor is wrong (or both?).
If the suspect 2-2) is guilty, it means we adjust PA current with very small range.
This is a completly safety way, but a wast of resource.
Now, the slider to control current adjustment indicate the output of DAC.
I will improve this to indicate current adjustment input, but it takes some time for me to learn about EPICS.
So after having broke Allegra by updating the kernel, I was able to get it running again by copying the xorg.conf.backup file over xorg.conf in /etc/X11. So at this point in time, Allegra is running with generic video drivers, as opposed to the ATI specific and proprietary drivers.
During the cleanup of the lab. Steve found a box with two BNCs going to the ICS DAQ interface and an unconnected D-SUB on the floor under the AP table. It seemed like a temperature sensor.
The BNCs were connected to C1:PEM-OSA_APTEMP and C1:PEM-OSA_SPTEMP.
Steve removed the box from the floor. These channels can be now used as spare DAQ channels. I labeled those cables.
So, near 2 of the trashcans in the control room and underneath a desk there are hundrends of ants. Is this normal?
After the ISS work, I aligned the IFO and confirmed that DRMI locks with good SPOB and AS166 values.
Yoichi and me found that the transfar function of the current shunt changed with the current of PA.
We changed PA current and fixed the unstability of ISS.
Now, laser power is stabilized finely, with band of about 1 Hz.
Yoich will post the stabilized noise spectrum.
There looks to be some non-linear relation between PA current and the TF of current shunt.
It had changed from the TF which we measured yesterday, so it might change again.
I try to write scripts to sweep PA current and measure the laser power and its rms automatically.
It will be apply for auto-adjustment of PA current.
Attached files are the transfar function of the current shunt with changing PA.
They have difference in lower frequency.
MOPAs and their settings, powers of 7 years in the 40m
I fixed the broken slider to change the current of the PA.
The problem was that the EPICS database assigned a wrong channel of the DAC to the slider.
I found that the PA current adjustment signal lines are connected to the CH3 &CH4 of VMIC4116 #1. However in the database file (/cvs/cds/caltech/target/c1psl/psl.db), the slider channel (C1:PSL-126MOPA_DCAMP) was assigned to CH2. I fixed the database file and rebooted c1psl. Then the PA current started to follow the slider value.
I moved the slider back and forth by +/-0.3V while the ISS loop was on. I observed that the amount of the low frequency fluctuation of the MOPA power changed with the slider position. At some current levels, the ISS instability problem went away.
Kakeru is now taking open-loop TFs and current shunt responses at different slider settings.
I measured the output noise of eache stage of ISS servo, and calcurated the noise ratio between input and
output of each stage.
Generaly, each noise ratio corresponds to their transfar function. This means servo filter works well, not
adding extra noise.
I attache example of them.
For 2nd stage, the noise ratio is smaller than transfar function with a few factor. This is because the
input noise is coverd by analyser's noise and ratio between output and input looks small.
This means the input noise of 2nd stage was enough small and all stage before 2nd stage work well